# DeCLIP: Decoding CLIP representations for deepfake localization

### v1

@@ 3.3 Detection with CLIP on Dolos
Importantly, the performance on local manipulations is not improved even if train on in-domain data: training on either P2 full or P2 local still yields only 79.2% (row 2) or 71.4% (row 4), respectively. On the other hand, Patch Forensics [3], which was used as a baseline in [59], is not affected local manipulations: it achieves an average precision of 95.3% (row 3). This result serves as our motivation for developing a patch-based approach on CLIP features. The full model, which we describe in the next section, is able to match the performance of Patch Forensics on local images, while maintaining good generalization performance.
@@ 4.2 Main results
Quantitative results. Our main results are shown in Table 2. For all methods we use all 16 train-test combinations of generators in Dolos and average IoU results for ID and OOD setups. Patch Forensics (row 1), the method that was originally applied for this task, shows good ID performance but performs poorly in OOD. Our starting point, the original CLIP method (row 3) has much worse performance in ID than Patch Forensics and comparable performance in OOD. Rows 6–8 show variants of our method, DeCLIP, which bring a significant boost to the original CLIP, improving both in ID and OOD setups. Compared to Patch Forensics, it has similar good performance in ID, but a 50% relative improvement in OOD. We also experiment with adding a larger decoder to Patch Forensics (row 2), but it did not help improve the OOD performance. Rows 4 and 5 show comparisons with other methods, PSCC-Net and CAT-Net that have been re-trained and tested in the exact same scenarios as DeCLIP. Both have significantly lower performance in ODD compared to DeCLIP. Their behaviour is different in ID: PSCC has good performance, while CAT-Net very poor.
Qualitative results. We show examples of output localization masks for all train-test setups produced by DeCLIP (ViT-L/14) and other methods in Figure 2. Notice that DeCLIP produced more consistent and clean masks even in the harder, OOD scenarios. PSCC-Net and Patch Forensics generally perform well in ID (with the exception of LDM–LDM case), but struggle in the OOD scenarios. CAT-Net and CLIP:ViT-L/14-linear seem to learn general face features, not related to the actual inpainted region.
@@ 4.3 Ablations
Backbones and representations depth. For both ViT-L/14 and ResNet-50 backbones we vary the depth of the layer at which we extract the pretrained representations. Our results are shown in Figure 3. When using ResNet-50, representations extracted at lower convolutional blocks (L1, L2) are best for manipulation localization in ID, while representations extracted at L3 block is best for OOD. The last block, L4 has lower performance both in ID and OOD. In the case of ViT-L/14, a similar trend can be seen with ID localization being higher when using features extracted at lower layers (L7) while OOD localization accuracy increasing when using higher level features (L21). Unlike ResNet-50, for ViT-L/14 there is no significant drop in performance at the last layer in the OOD scenario.
@@ 5 The case of LDM-inpainted images
Larger models improve performance on LDM, but capacity alone is not sufficient for generalization. Patch Forensics fails at localizing even ID manipulations on LDM images, while DeCLIP works better. An important difference between the two is the larger decoder employed by the latter. We verify whether the reason for the difference in performance is solely based on network capacity. We consider two larger variants: Patch Forensics but with the conv-20 decoder (42.6M parameters) and PSCC [36] (3.6M parameters); both of these networks are trained from scratch on the LDM subset. The results in Table 5 show that indeed the network capacity is responsible to a degree for the good performance, as the larger variant of Patch Forensics improves substantially over the smaller baseline. However, PSCC is better than both Patch Forensics variants, while DeCLIP achieves the best performance, at a number of parameters comparable to the larger Patch Foresnics variant.
@@
Table 6: Comparison in terms of IoU to other methods on the more general domain provided by MS COCO. DeCLIP shows a more stable performance across all datasets. Values in bold indicate best results in each of the two sections.
@@
Conclusions translate to more general domains. While Dolos dataset enabled a careful analysis of deepfake localization in a challenging and realistic setting, it covers only a narrow domain: faces. We verify whether the main conclusions apply to more general images. We inpaint almost 11k images (9k train, 829 validation, 985 test) from MS COCO [33] with Stable Difusion [53], which is also an LDM model. We select the mask of a random object whose area is larger than 5% and prompt the inpainting model with the original image caption. This dataset, which we name COCO-SD, is used for training the localization models. To evaluate generalization, we use the AutoSplice dataset [24]. This dataset consists of images manipulated by DALLE-2 [50], where specific objects are replaced with other objects. It has three variants differing in terms of the JPEG compression applied: the original uncompressed one (quality factor: 100) and two compressed ones (quality factors of 75 and 90). Table 6 presents the results for DeCLIP, Patch Foresnics, PSCC and CAT-Net. We also show results of pretrained models, for the two last ones, MantraNet [66] and TruFor [15]. We observe that the pretrained models, especially CAT-Net and PSCC, are unstable: they have good performance on some of the datasets, but equally poor on others. MantraNet performs poorly on all datasets, while TruFor has good performance on AutoSplice-100 but lower on its JPEG compressed variants and COCO-SD. Instead, we see that training on the LDM-based COCO-SD dataset offers more stable performance. This is especially true for Patch Forensics, PSCC and DeCLIP, which perform similarly across all three AutoSplice variants. In terms of model comparison, DeCLIP works better in-domain, while slightly outperforming PSCC on the out-of-domain AutoSplice dataset. The visual examples in Figure 6 show that DeCLIP produces more precise localization of the forged object.
Figure 6: Manipulation localization results on COCO-SD, which has a more challenging set of masks and diverse content. DeCLIP offers a more precise localization of the manipulated area.
@@ 6 Conclusion
Our paper presents DeCLIP—a first attempt at decoding large self-supervised representations for manipulation localization. Through extensive experiments we showed that, not only is manipulation localization feasible using these features, but they also significantly improve generalization capabilties in OOD scenarios, with train–test generator mismatch. We conducted a comprehensive analysis of the factors that contribute to the successful decodings of those features: backbone type, layer depth, decoder type and size. Our findings reveal that larger, convolutional decoders improve the quality of the predicted masks compared to linear or self-attention ones. Moreover, VIT-L/14 and ResNet-50 backbones show contrasting behavior when looking at detailed train-test scenarios that can be exploited by combining representations from both backbones. Finally, we showed that, contrary to prior assumptions, manipulation localization can be effectively performed even in the challenging case of LDM data. Interestingly, learning on this type of data offers robustness and improves generalization to other types of local manipulations.
Acknowledgements. This work was funded by EU Horizon projects AI4TRUST (No. 101070190) and ELIAS (No. 101120237).

### v2

@@ 3.3 Detection with CLIP on Dolos
Importantly, the performance on local manipulations is not improved even if we train on in-domain data: training on either P2 full or P2 local still yields only 79.2% (row 2) or 71.4% (row 4), respectively. On the other hand, Patch Forensics [3], which was used as a baseline in [60], is not affected local manipulations: it achieves an average precision of 95.3% (row 3). This result serves as our motivation for developing a patch-based approach on CLIP features. The full model, which we describe in the next section, is able to match the performance of Patch Forensics on local images, while maintaining good generalization performance.
@@ 4.2 Main results
Quantitative results. Our main results are shown in Table 2. For all methods we use all 16 train-test combinations of generators in Dolos and average IoU results for ID and OOD setups. Patch Forensics (row 1), the method that was originally applied for this task, shows good ID performance but performs poorly in OOD. Our starting point, the original CLIP method (row 3) has much worse performance in ID than Patch Forensics and comparable performance in OOD. Rows 6–8 show variants of our method, DeCLIP, which bring a significant boost to the original CLIP, improving both in ID and OOD setups. Compared to Patch Forensics, it has similar good performance in ID, but a 50% relative improvement in OOD. We also experiment with adding a larger decoder to Patch Forensics (row 2), but it did not help improve the OOD performance. Rows 4 and 5 show comparisons with other methods, PSCC-Net and CAT-Net that have been re-trained and tested in the exact same scenarios as DeCLIP. Both have significantly lower performance in OOD compared to DeCLIP. Their behaviour is different in ID: PSCC has good performance, while CAT-Net very poor.
Qualitative results. We show examples of output localization masks for all train–test setups produced by DeCLIP (ViT-L/14) and other methods in Figure 2. Notice that DeCLIP produces more consistent and clean masks even in the harder, OOD scenarios. PSCC-Net and Patch Forensics generally perform well in ID (with the exception of LDM–LDM case), but struggle in the OOD scenarios. CAT-Net and CLIP:ViT-L/14-linear seem to learn general face features, not related to the actual inpainted region.
@@ 4.3 Ablations
Backbones and representations depth. For both ViT-L/14 and ResNet-50 backbones we vary the depth of the layer at which we extract the pretrained representations. Our results are shown in Figure 3. When using ResNet-50, representations extracted at lower convolutional blocks (L1, L2) are best for manipulation localization in ID, while representations extracted at L3 block is best for OOD. The last block, L4 has lower performance both in ID and OOD. In the case of ViT-L/14, a similar trend can be seen with ID localization being higher when using features extracted at lower layers (L7), while OOD localization accuracy increasing when using higher level features (L21). Unlike ResNet-50, for ViT-L/14 there is no significant drop in performance at the last layer in the OOD scenario.
@@ 5 The case of LDM-inpainted images
Larger models improve performance on LDM, but capacity alone is not sufficient for generalization. Patch Forensics fails at localizing even ID manipulations on LDM images, while DeCLIP works better. An important difference between the two is the larger decoder employed by the latter. We verify whether the reason for the difference in performance is solely based on network capacity. We consider two larger variants: Patch Forensics but with the conv-20 decoder (42.6M parameters) and PSCC [36] (3.6M parameters); both of these networks are trained from scratch on the LDM subset. The results in Table 5 show that indeed the network capacity is responsible to a degree for the good performance, as the larger variant of Patch Forensics improves substantially over the smaller baseline. However, PSCC is better than both Patch Forensics variants, while DeCLIP achieves the best performance, at a number of parameters comparable to the larger Patch Forensics variant.
@@
Table 6: Localization performance (in terms of IoU) on the general domain of MS COCO. DeCLIP shows stable performance across all datasets. Bold indicates best results in each of the sections.
@@
Conclusions translate to more general domains. While the Dolos dataset enabled a careful analysis of deepfake localization in a challenging and realistic setting, it covers only a narrow domain: faces. We verify whether the main conclusions apply to more general images. We inpaint almost 11k images (9k train, 829 validation, 985 test) from MS COCO [33] with Stable Diffusion [54], which is also an LDM model. We select the mask of a random object whose area is larger than 5% and prompt the inpainting model with the original image caption. This dataset, which we name COCO-SD, is used for training the localization models. (A similar dataset has been proposed by concurrent work [40].) To evaluate generalization, we use the AutoSplice dataset [24]. This dataset consists of images manipulated by DALLE-2 [51], where specific objects are replaced with other objects. It has three variants differing in terms of JPEG compression: the original uncompressed one (quality factor: 100) and two compressed ones (quality factors of 75 and 90). Table 6 presents the results for DeCLIP, Patch Forensics, PSCC and CAT-Net. We also show results of pretrained models for the two last methods, as well as for MantraNet [67] and TruFor [15]. We observe that the pretrained models, especially CAT-Net and PSCC, are unstable: they have good performance on some of the datasets, but equally poor on others. MantraNet performs poorly on all datasets, while TruFor has good performance on AutoSplice-100 but lower on its JPEG compressed variants and COCO-SD. Instead, training on the LDM-based COCO-SD dataset offers more stable performance. This is especially true for Patch Forensics, PSCC and DeCLIP, which perform similarly across all three AutoSplice variants. In terms of model comparison, DeCLIP works better in-domain, while slightly outperforming PSCC on the out-of-domain AutoSplice dataset. The visual examples in Figure 6 show that DeCLIP produces more precise localization of the forged object.
Figure 6: Manipulation localization on COCO-SD. DeCLIP offers a more precise localization than the other methods.
@@ 6 Conclusion
Our paper presented DeCLIP—a first attempt at decoding large self-supervised representations for manipulation localization. Through extensive experiments we showed that, not only is manipulation localization feasible using these features, but they also significantly improve generalization capabilities in OOD scenarios, when there is a train–test generator mismatch. We conducted a comprehensive analysis of the factors that contribute to the successful decoding of those features: backbone type, layer depth, decoder type and size. We found that larger, convolutional decoders improve the quality of the predicted masks compared to linear or self-attention ones. Moreover, VIT-L/14 and ResNet-50 backbones show contrasting behavior that can be exploited by combining representations from both backbones. Finally, we showed that, contrary to prior assumptions, manipulation localization can be effectively performed even in the challenging case of LDM data. Interestingly, learning on this type of data offers robustness and improves generalization to other types of local manipulations.
Acknowledgements. This work was supported in part by the EU Horizon projects AI4TRUST (No. 101070190) and ELIAS (No. 101120237), and by CNCS-UEFISCDI (PN-IV-P7-7.1-PTE-2024-0600).